KTeV Shift Summary for DAY 9/27/99

S. Elizabeth Turner (eturner@kudzu)
Mon, 27 Sep 1999 15:27:29 -0500

Shift Summary for DAY shift, 27/SEP/99

On Shift: H. White, E. Turner

Hours of beam delivered at all intensities: 8.0/8.5
Hours of beam delivered at nominal intensity: 8.0/8.5
Hours of E799 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

Various problems with tapes/DA init-ing run 14744.
Switchover to special run configuration w/ no absorber.
========================================================================
Runs ended this shift:

14737 BEAM_799 (16.8 M events)
14744 BEAM_799 (4.7 M events)

Tapes brought to FCC this shift:

UP*034-036, UPS012
========================================================================

Accesses during shift: none

========================================================================
Overview of shift: 799 datataking, special 799 run w/ no absorber

========================================================================
Description of Shift:

7:15 Run 14737 ends -- tapes full. We change tapes. New tapes are
UP*037-039,UPS013.

7:20 to 9:00 Many problems getting the run to start. Clean ktev3 drive 1.
Clean ktev1 drive 2. Replace tape in ktev1 drive 2 with new
tape -- suspect old tape is bad. Reset pipeline several times,
eventually reboot pipeline crate 2. Log in and out of da_ktev3
session which runs run control. ~5 drt_fresh_starts during
all this.

9:00 14744 starts successfully with 2 kfrends and 2 pipe resets :)

10:57 MCR tells us that beam intensity will be lowered to reset QXR.

11:45 Stop run 14744 to switch to no absorber/lower intensity mode.

12:06 Start run 14745 -- special 799 run with no absorber

12:45 Note that 'weak' channel in RC6 is due to ped change. Update ADGN
constants before next run.

13:21 We request that MCR keep 5.5E12 as upper limit on intensity.

14:27 Some small problems found in online kumacs (target#2V, trg#2V,
TRD ADC avg) see XXII.3 for details.
========================================================================
Follow up on Notes from Previous Shift

========================================================================
Items to Follow up on Next Shift

Switch back to absorber-in configuration and return to high intensity.
Update ADGN constants before next run.